# QCircuitNet: A Large-Scale Hierarchical Dataset for Quantum Algorithm Design

### v1

@@ 5 Experiments
5.1 Benchmarking LLMs on QCircuitNet
We benchmark the quantum algorithm design capabilities of leading closed-source and open-source large language models using QCircuitNet. The workflow of our benchmark is illustrated in Figure 2. The total computation cost is approximately equivalent to two days on an A100 GPU.
Figure 2: Flowchart of benchmarking QCircuitNet.
@@ 5.1 Benchmarking LLMs on QCircuitNet
Recently, the GPT series models have become the benchmark for generative models due to their exceptional performance. Specifically, we include two models from OpenAI, GPT-3.5-turbo (Brown et al., 2020) and GPT-4 (OpenAI et al., 2024), in our benchmark. Additionally, the LLAMA series models (Touvron et al., 2023a; b) are widely recognized as leading open-source models, and we have selected LLAMA-3-8B for our study. For a comprehensive evaluation, we also benchmark Phi-3-medium-128k (Abdin et al., 2024) and Mistral-7B-v0.3 (Jiang et al., 2023).
Prompts.
We employ a few-shot learning framework, a prompting technique that has shown considerable success in generative AI (Xie et al., 2021). In this approach, we utilize either 1 or 5 examples, followed by a problem description. To ensure we do not train and test on the same quantum algorithm, we implement k-fold validation. This method involves using one problem as the test set while the remaining problems serve as the training set, rotating through each problem one at a time.
@@
We use three evaluation metrics (see Appendix C.1 for more details):
@@
    BLEU Score: this metric measures how closely the generated code matches the reference code, with a higher BLEU score indicating greater similarity.
@@
    Verification function: this function checks the syntax validation and the result correctness of the code produced by the language model.
@@
    Byte Perplexity: this metric evaluates the model’s ability to predict the next byte in a sequence. Lower byte perplexity indicates better performance by reflecting the model’s predictive accuracy.
The results for BLEU and verification function score are shown in Figure 3, Table 1, and Table 2. We include the results of Byte Perplexity in Appendix C.2.
Figure 3: Benchmarking algorithm design and oracle construction in BLEU scores.
Table 1: Benchmarking algorithm design in verification function scores.
@@
Avg
gpt4o 1 -0.8462 -0.5538 -0.7089 -0.8423 -1.$0$0.6692 -0.8462 -1.$0$1.$0$1.$0$0.8844 -0.6667 -0.8348
(±0.1042) (±0.1986) (±0.1879) (±0.0000) (±0.0000) (±0.1447) (±0.1538) (±0.0000) (±0.0000) (±0.0000) (±0.1156) (±0.3333)
gpt4o 5 -0.3054 0.0135 -0.2071 -0.5846 -0.6154 -0.3692 -0.1538 -0.4967 -0.8700 -0.9231 -0.4889 $0$0.4167
(±0.2086) (±0.2070) (±0.2089) (±0.3107) (±0.1804) (±0.1443) (±0.2738) (±0.2210) (±0.1300) (±0.0769) (±0.2077) (±0.0000)
Llama3 1 -0.2308 -0.7692 -0.7143 -1.$0$0.9231 -1.$0$0.6154 -0.9285 -1.$0$0.3846 -1.$0$1.$0$0.7972
(±0.2571) (±0.1216) (±0.1844) (±0.1429) (±0.0769) (±0.0000) (±0.1804) (±0.0715) (±0.0000) (±0.1404) (±0.0000) (±0.0000)
Llama3 5 0.0769 -0.2308 -0.5393 -0.9231 -0.7692 -0.8462 -0.3846 -0.7276 -1.$0$0.1538 -0.8889 -1.$0$0.6155
(±0.2107) (±0.1662) (±0.2185) (±0.0000) (±0.1662) (±0.1042) (±0.1404) (±0.1468) (±0.0000) (±0.1042) (±0.1111) (±0.0000)
gpt3.5 1 -0.8462 -0.7154 -0.5679 -1.$0$1.$0$0.6231 -0.8462 -1.$0$1.$0$1.$0$0.7778 -1.$0$0.8647
(±0.1042) (±0.1503) (±0.2037) (±0.0000) (±0.0000) (±0.1680) (±0.1538) (±0.0000) (±0.0000) (±0.0000) (±0.1470) (±0.0000)
gpt3.5 5 -0.6154 -0.0571 -$0$0.9154 -0.6538 -0.1646 -0.2308 -0.4513 -0.8778 -0.8462 -0.3311 $0$0.4328
(±0.1404) (±0.1934) (±0.1687) (±0.0000) (±0.1538) (±0.1395) (±0.2809) (±0.2410) (±0.1222) (±0.1042) (±0.1672) (±0.0000)
Phi3 1 -0.8462 -0.7750 -1.$0$1.$0$1.$0$1.$0$0.3846 -1.$0$0.8878 -0.8462 -1.$0$1.$0$0.8950
(±0.1538) (±0.1527) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.2130) (±0.0000) (±0.1122) (±0.1042) (±0.0000) (±0.0000)
Phi3 5 -0.6577 -0.3821 -0.8286 -0.6923 -1.$0$0.6100 -0.9231 -0.3569 -0.8333 -0.8462 -0.8889 -1.$0$0.7516
(±0.1891) (±0.2342) (±0.1714) (±0.0000) (±0.0000) (±0.1425) (±0.0769) (±0.2402) (±0.1667) (±0.1042) (±0.1111) (±0.0000)
Mistral 1 -0.8462 -0.8590 -0.7107 -1.$0$1.$0$0.9192 -0.7692 -1.$0$1.$0$0.6923 -0.8889 -1.$0$0.8905
(±0.1042) (±0.1410) (±0.1868) (±0.0000) (±0.0000) (±0.0808) (±0.1662) (±0.0000) (±0.0000) (±0.1332) (±0.1111) (±0.0000)
Mistral 5 -0.6246 -0.6667 -0.4071 -1.$0$0.9231 -0.9115 -0.6923 -0.8820 -1.$0$0.5385 -0.8889 -0.6667 -0.7668
(±0.1664) (±0.1820) (±0.2106) (±0.1429) (±0.0769) (±0.0885) (±0.1332) (±0.1180) (±0.0000) (±0.1439) (±0.1111) (±0.3333)
Table 2: Benchmarking oracle construction in verification function scores.
Model Shot Bernstein-Vazirani Deutsch-Jozsa Diffusion-Operator Grover Simon
Generalized-
Simon (multi-str)
Avg
gpt4o 1 -0.3200 -$0$0.8462 -0.9885 -0.4674 -0.3750 -0.5012
(±0.0530) (±0.0438) (±0.1538) (±0.0115) (±0.0545) (±0.0870)
gpt4o 5 -0.1100 $0$0.3077 -0.9540 -0.0870 -0.3125 -0.2819
(±0.0399) (±0.0506) (±0.2083) (±0.0279) (±0.0295) (±0.0832)
Llama3 1 -0.7300 -$0$0.3846 -1.$0$0.6848 -0.9375 -0.7061
(±0.0468) (±0.0704) (±0.1404) (±0.0000) (±0.0487) (±0.0435)
Llama3 5 -0.0500 0.1700 -0.8462 -1.$0$0.6413 -0.6875 -0.5092
(±0.0359) (±0.0551) (±0.1042) (±0.0000) (±0.0503) (±0.0832)
gpt3.5 1 -0.3500 -$0$0.8462 -1.$0$0.3696 -1.$0$0.6010
(±0.0539) (±0.0470) (±0.1538) (±0.0000) (±0.0529) (±0.0000)
gpt3.5 5 -0.1100 $0$0.3077 -0.9770 -0.1087 -0.4063 -0.3149
(±0.0373) (±0.0531) (±0.2083) (±0.0162) (±0.0326) (±0.0989)
Phi3 1 -0.6800 -0.6100 -0.9231 -1.$0$0.7500 -1.$0$0.8272
(±0.0510) (±0.0584) (±0.0769) (±0.0000) (±0.0454) (±0.0000)
Phi3 5 -0.5400 -0.4300 -1.$0$1.$0$0.8370 -0.9063 -0.7855
(±0.0521) (±0.0685) (±0.0000) (±0.0000) (±0.0387) (±0.0524)
Mistral 1 -$0$0.4300 -0.9231 -0.9540 -0.6087 -0.9375 -0.7089
(±0.0512) (±0.0640) (±0.0769) (±0.0279) (±0.0512) (±0.0435)
Mistral 5 -0.3700 -0.1300 -1.$0$0.9195 -0.2391 -0.9063 -0.5942
(±0.0506) (±0.0734) (±0.0000) (±0.0373) (±0.0447) (±0.0524)
The results illustrate that most models achieve better scores in the five-shot setting, which indicates their ability to learn effectively from contextual examples. Notably, models perform well on tasks like Bernstein-Vazirani and Deutsch-Jozsa but struggle with more complex algorithms such as Grover, phase estimation, and quantum Fourier transform, highlighting differences in task difficulty. Furthermore, although the BLEU scores show a general trend of consistency with verification scores, some discrepancies arise, such as the swap test showing relatively high BLEU scores but incorrect algorithm generation by most models. This observation emphasizes the need for complementary evaluation metrics beyond BLEU to accurately assess model performance, which highlights the importance of our verification function. Additionally, GPT-4o and GPT-3.5 consistently excel in long-context comprehension, significantly outperforming other models across tasks, which highlights their superior in-context learning capabilities. For more detailed analysis of the experimental results, we refer to Appendix C.1.
@@
In Appendix C.3, we include several case studies to illustrate and analyze various types of errors made by LLMs. For example, GPT-4o tends to use advanced OpenQASM 3.0 features unsupported by Qiskit yet and novel namespace which might result in global conflicts in one-shot setting. This tendency to improvise by drawing on pre-trained knowledge rather than closely following the syntax of the example leads to avoidable "errors" and low verification scores. This issue is significantly alleviated in the 5-shot setting, highlighting GPT-4o’s strong in-context learning ability.
5.2 Fine-tuning on QCircuitNet
Although QCircuitNet is targeted as a benchmark dataset at current stage, we consider fine-tuning / training from scratch based on our dataset as an interesting and important research direction. The unique nature of quantum data requires novel fine-tuning methods and model architecture designs, which could serve as a standalone topic. For a primitive demonstration, we present fine-tuning results on data from the oracle construction task here.
Following Dettmers et al. (2024), we quantize the model to 8-bits and then train it with LORA (Hu et al., 2022). In our experiments, we use fp16 computational datatype. We set LoRA $r=16,\alpha=32$ and add LoRA modules on all the query and value layers. We also use AdamW (Loshchilov & Hutter, 2019) and LoRA dropout of $0.05$. The results are shown as follows:
Table 3: Fine-tuning oracle construction scores.
Score Model Setting
Bernstein-
Vazirani
Deutsch-
Jozsa
Grover Simon Clifford Universal Avg

### v2

@@ 4 Experiments
4.1 Benchmarking LLMs on QCircuitBench
We benchmark the quantum algorithm design capabilities of leading closed-source and open-source large language models using QCircuitBench. The workflow of our benchmark is illustrated in Figure 2. The total compute footprint per model includes $\approx$48 A100-GPU hours for inference and an additional 24 single-core CPU hours for verification and simulation.
Figure 2: Flowchart of benchmarking QCircuitBench.
@@ 4.1 Benchmarking LLMs on QCircuitBench
We evaluate a range of state-of-the-art LLMs, including both proprietary and open-source representatives. From OpenAI, we include GPT-4o (openai2024gpt4), which serve as widely adopted baseline due to its strong general-purpose performance. For open-source models, we benchmark LLAMA-3-8B (meta2024llama3), a prominent member of the LLAMA family (touvron2023llama; touvron2023llama2; meta2024llama3) known for its efficiency and competitive performance. We further include Qwen 2.5 (qwen2.5) and DeepSeek-R1 (deepseekr1), two recently released models that have shown promising results across a variety of reasoning and code generation tasks. To contextualize model performance, we conduct a human study and report human baseline scores at the bottom line of each table as reference. The instruction for human study is presented in Appendix A.3.
All models are evaluated using a few-shot prompting strategy, a widely used method in generative language model evaluation (brown2020language; xie2021explanation; dong-etal-2024-survey). We provide 1 or 5 in-context examples, followed by the task description. To prevent data leakage, we use $k$-fold cross validation, holding out each quantum algorithm from the prompt during evaluation.
@@
We use four evaluation metrics:
@@
    BLEU Score: This metric measures how closely the generated code matches the reference code, with a higher BLEU score indicating more similarity.
@@
    Byte Perplexity: Evaluates the model’s ability to predict the next byte in a sequence. Lower PPL suggests stronger autoregressive modeling and more coherent outputs.
@@
    Verification Function: This function performs syntax checking and semantic evaluation of the code produced by language models.
    To be specific, we evaluate the result using three criteria:
    1. (a)
      QASM Syntax: We first check the syntax of the OpenQASM code provided by the model. The syntax verification function $V_{\text{QASM}}(q)$ is set to be 1 if the QASM syntax is correct, and 0 otherwise.
    2. (b)
      Python Syntax: Similarly, the syntax of the post-processing Python code, denoted $V_{\text{code}}(c)$, is set to be 1 if the Python syntax is correct, and 0 otherwise.
    3. (c)
      Semantic: If both syntax checks pass, we proceed to evaluate the functional correctness. For quantum algorithm design and oracle construction tasks, for each test case $t$, we run the quantum circuit simulation for a number of shots $M$, and compare the result with the ground truth. The score reflects the success rate over test cases, which is calculated as:
      $E_{\text{sem}}=\frac{\sum_{t=1}^{T}\sum_{m=1}^{M}\mathbb{I}[\text{result = ground-truth}]}{T\times M}.$
      For state preparation and random circuit synthesis tasks, let $|\psi^{\star}\rangle$ be the target state and $|\psi\rangle$ be the state produced by the model’s circuit, the semantic score is defined as the fidelity between the generated and target quantum states:
      $E_{\text{sem}}=|\langle\psi^{\star}|\psi\rangle|^{2},$
      Overall, this task-specific semantic score $E_{\text{sem}}$ provides a principled and quantitative measure of quantum correctness tailored to each task.
    The final verification score is a triplet $(V_{\text{QASM}}(q),V_{\text{code}}(c),E_{\text{sem}})$.
  4. 4.
    Efficiency Evaluation: We incorporate another three metrics to evaluate the efficiency of generated algorithms.
    1. (a)
      Gate Count Ratio: the number of quantum gates used by the model, divided by that of the reference implementation. A lower ratio indicates higher gate-level efficiency.
    2. (b)
      Shot Count Ratio: the query complexity required by the model, divided by the reference. Again, a smaller value indicates better efficiency.
    3. (c)
      Time Count Ratio: the total execution time of the model, divided by the reference. A lower ratio reflects better runtime performance.
The BLEU scores of all three tasks are shown in Figure 3. The qasm syntax and semantic verification scores of quantum algorithm design task are shown in Table 1 and Table 2. We include the results of python syntax verification, efficiency metrics, Byte Perplexity, and open-book setting of quantum algorithm design task in Appendix LABEL:appendix:quantum-algorithm-design. Other results for oracle construction and random circuit synthesis tasks are included in Appendix LABEL:appendix:benchmark-oracle-random. Benchmarking results for the Cirq implementation are presented in Appendix LABEL:appendix-cirq.
Figure 3: Benchmarking algorithm design and oracle construction tasks in BLEU scores.
Table 1: QASM syntax score for benchmarking quantum algorithm design.
@@
Shor VQE QAOA QAE ENC Avg
GPT-4o 1 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 1.0000 1.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.2308 1.0000 0.8333 0.5833 0.2734
(±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0843) (±0.0000) (±0.0904) (±0.1486)
GPT-4o 5 1.0000 1.0000 0.0000 0.6154 0.5385 0.9231 0.5714 1.0000 1.0000 0.4444 0.0769 0.1111 0.0000 0.2308 0.7222 1.0000 0.5833 0.5775
(±0.0000) (±0.0000) (±0.0000) (±0.1404) (±0.1439) (±0.0769) (±0.2020) (±0.0000) (±0.0000) (±0.1757) (±0.0769) (±0.1111) (±0.0000) (±0.0843) (±0.1086) (±0.0000) (±0.1486)
Llama3 1 0.1538 0.2308 0.3077 0.4615 0.0000 0.1538 0.1429 0.4615 0.1429 0.3333 0.5385 0.4444 0.0000 0.2574 0.1667 0.0000 0.3438 0.2435
(±0.1042) (±0.1216) (±0.1332) (±0.1439) (±0.0000) (±0.1042) (±0.1429) (±0.1439) (±0.0971) (±0.1667) (±0.1439) (±0.1757) (±0.0000) (±0.0285) (±0.0544) (±0.0000) (±0.0853)
Llama3 5 0.5385 0.3846 0.6154 0.5385 0.3846 0.1538 0.2857 0.9231 0.5000 0.3333 0.8462 0.3333 0.0000 0.2363 0.9375 0.0000 0.8125 0.4602
(±0.1439) (±0.1404) (±0.1404) (±0.1439) (±0.1404) (±0.1042) (±0.1844) (±0.0769) (±0.1387) (±0.1667) (±0.1042) (±0.1667) (±0.0000) (±0.0277) (±0.0353) (±0.0000) (±0.0701)
Qwen 2.5 1 0.0769 0.1538 0.0000 0.0769 0.0769 0.3077 0.4286 0.2308 0.2857 0.2222 0.5385 0.1111 0.0000 0.4515 0.8750 0.0000 1.0000 0.2844
(±0.0769) (±0.1042) (±0.0000) (±0.0769) (±0.0769) (±0.1332) (±0.2020) (±0.1216) (±0.1253) (±0.1470) (±0.1439) (±0.1111) (±0.0000) (±0.0324) (±0.0482) (±0.0000) (±0.0000)
Qwen 2.5 5 0.3077 0.6154 0.1538 0.3077 0.2308 0.1538 0.4286 0.6154 0.5714 0.2222 0.4615 0.2222 0.0000 0.3544 0.9583 1.0000 0.7188 0.4307
(±0.1332) (±0.1404) (±0.1042) (±0.1332) (±0.1216) (±0.1042) (±0.2020) (±0.1404) (±0.1373) (±0.1470) (±0.1439) (±0.1470) (±0.0000) (±0.0311) (±0.0291) (±0.0000) (±0.0808)
DeepSeek-R1 1 0.0000 0.0769 0.0000 0.0000 0.0000 0.0000 0.1429 0.0769 0.0714 0.0000 0.1538 0.0000 0.0000 0.07173 0.2292 0.0000 0.1563 0.0576
(±0.0000) (±0.0769) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.1429) (±0.0769) (±0.0714) (±0.0000) (±0.1042) (±0.0000) (±0.0000) (±0.0168) (±0.0613) (±0.0000) (±0.0652)
DeepSeek-R1 5 0.3846 0.0769 0.0000 0.0769 0.0769 0.0000 0.0000 0.1538 0.1429 0.0000 0.2308 0.0000 0.0000 0.0084 0.4167 1.0000 0.4375 0.1768
(±0.1404) (±0.0769) (±0.0000) (±0.0769) (±0.0769) (±0.0000) (±0.0000) (±0.1042) (±0.0971) (±0.0000) (±0.1216) (±0.0000) (±0.0000) (±0.0060) (±0.0719) (±0.0000) (±0.0891)
Human - 0.5000 1.0000 0.0000 1.0000 1.0000 1.0000 1.0000 0.0000 0.0000 1.0000 1.0000 0.0000 0.5000 1.0000 1.0000 1.0000 0.6667 0.6862
Table 2: Semantic score for benchmarking quantum algorithm design.
Model Shot
Bernstein
Vazirani
Deutsch
Jozsa
Grover
Phase
Estimation
QFT Simon GHZ
Random
Number
Generator
Swap
Test
W
State
Generalized
Simon
(multi-str)
Generalized
Simon
Shor VQE QAOA QAE ENC Avg
GPT-4o 1 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.4167 0.0245
(±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.1205)
GPT-4o 5 1.0000 1.0000 0.0000 0.0846 0.0000 0.0923 0.0000 0.0000 0.7852 0.1156 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.3333 0.2006
(±0.0000) (±0.0000) (±0.0000) (±0.0576) (±0.0000) (±0.0769) (±0.0000) (±0.0000) (±0.0203) (±0.0585) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.1421)
Llama3 1 0.0000 0.0154 0.0019 0.0000 0.0000 0.0000 0.0000 0.0000 0.0591 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.1101 0.0110
(±0.0000) (±0.0154) (±0.0019) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0591) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0526)
Llama3 5 0.0442 0.2269 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0986 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.1799 0.0323
(±0.0442) (±0.1017) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0673) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0482)
Qwen 2.5 1 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
(±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000)
Qwen 2.5 5 0.0154 0.2854 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0042 0.1042 1.0000 0.2881 0.0998
(±0.0154) (±0.0947) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0042) (±0.0446) (±0.0000) (±0.0801)
DeepSeek-R1 1 0.0000 0.0600 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0605 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0938 0.0126
(±0.0000) (±0.0600) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0605) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0524)
DeepSeek-R1 5 0.0788 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0948 0.0102
(±0.0768) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0000) (±0.0387)
Human - 0.2500 0.0000 0.0000 0.8500 0.0000 0.0000 0.0000 0.0000 0.0000 0.5000 0.5000 0.0000 0.0000 0.0000 0.0000 0.2357 0.0000 0.1374
We observe the following phenomena from the results:
  • We observe notable discrepancies among BLEU, syntax, and semantic scores. While BLEU loosely correlates with verification outcomes, counter-examples exist. For instance, the GHZ state attains relatively high BLEU but low semantic accuracy. Although models achieve relatively high QASM syntax scores on several algorithms, their overall semantic performance remains poor. This observation highlights the necessity of our multi-faceted verification function.
  • Few-shot learning improves model performance on simpler tasks, demonstrating the models’ ability to learn effectively from contextual examples. Notably, in the Bernstein-Vazirani Algorithm, the QASM syntax score of GPT-4o improves from 0.0000 to 1.0000, and in the Deutsch-Jozsa Algorithm, Qwen 2.5 achieves an increase in semantic score from 0.0000 to 0.2854. However, even with few-shot prompting, all models struggle on more complex algorithms such as Quantum Fourier Transform, Generalized Simon’s Problem, and Shor’s Algorithm, revealing clear differences in task difficulty.
  • Near-term quantum algorithms pose persistent challenges. In particular, for variational quantum algorithms, models frequently fail to construct correct parameterized circuits or apply optimization procedures appropriately. For example, DeepSeek-R1 achieves all zero scores on semantic score of VQE and QAOA, reflecting the current limitations of LLMs in modeling hybrid quantum-classical workflows.
  • GPT-4o consistently excels at long-context comprehension, outperforming other models across tasks and demonstrating superior in-context learning. In contrast, DeepSeek-R1 often underperforms due to its long-chain reasoning style, which often exceeds the context window before producing a complete and verifiable solution.
@@
In Appendix LABEL:sec:case-study, we include several case studies to illustrate and analyze various types of errors made by LLMs. In particular, they can be summarized as follows:
  • Improvisation error: GPT-4o tends to use advanced OpenQASM 3.0 features unsupported by Qiskit yet and novel namespace which might result in global conflicts in one-shot setting. This tendency to improvise by drawing on pre-training knowledge rather than closely following the syntax of the example leads to avoidable "errors" and low verification scores. This issue is significantly alleviated in the 5-shot setting, highlighting GPT-4o’s strong in-context learning ability. A detailed case study is given in Appendix LABEL:sec:Improvisation.
  • Counting error: LLMs often fail to correctly identify the positions of ‘1’s in a binary string when constructing oracles for problems such as Bernstein-Vazirani. For instance, given the secret string $s=000101$, GPT-4o misplaces the control qubits for CX gates, despite being explicitly reminded of the correct rule and asked to list the indices with value 1. This misidentification highlights a fundamental limitation of LLMs in performing basic indexing tasks. A detailed case study is given in Appendix C.3.2.
  • Data contamination: We observe a performance separation between writing general Qiskit codes and explicit gate-level circuits in QASM. Since Qiskit provides detailed tutorial with general codes for several algorithms, LLMs may rely on memorization and retrieval rather than genuine algorithm design. Our dataset, based on QASM files created from scratch, may help circumvent this issue and serve as a stable and fair method for benchmarking AI syntax learning. A detailed case study is given in Appendix C.3.3.
4.2 Fine-tuning on QCircuitBench
Although QCircuitBench is targeted as a benchmark dataset at current stage, we consider fine-tuning or training from scratch based on our dataset as an interesting and important research direction. The unique nature of quantum data requires novel fine-tuning methods and model architecture designs, which could serve as a standalone topic. As a preliminary demonstration, we present the results of fine-tuning the Llama3-8B model on the oracle construction and random circuit synthesis tasks here.
Following the QLoRA approach by dettmers2024qlora, we first quantize the model to 8-bit precision and then fine-tune it using LoRA (hu2022lora). Our experiments use the fp16 data type and apply LoRA with rank $r=16$, scaling factor $\alpha=32$, and modules inserted into all query and value projection layers. We adopt the AdamW optimizer (loshchilov2018decoupled) and set the LoRA dropout rate to $0.05$. We define an aggregated verification score to represent the overall syntax and semantic performance of the model. If grammar errors exist, the function returns -1; if the program can execute successfully, the function returns a score between [0, 1] indicating the success rate on test cases. The results are shown as follows:
Figure 4: Fine-tuning oracle construction scores.
Score Model Setting Bernstein-Vazirani Deutsch-Jozsa Grover Simon Clifford Universal Avg